# Unified Evolution of Electromagnetic Sources in Homogeneous Fields

### v1

@@ I Introduction and Summary
For particles with spin in very strong or inhomogeneous fields additional terms arise, which originate from a Stern-Gerlach-type force. A covariant formulation of the Stern-Gerlach force was first considered by Frenkel [3], then by Good [26] and Nyborg [27], among others. It has been revisited and extended in recent years [28], and the dynamics of both neutral and charged particles were studied in [29, 30, 31]. We do not consider such contributions here since we focus on homogeneous fields and fields, whose gradients are negligibly small. Further extensions to the dynamical equations and implications thereof are presented in [32]. In the framework of general relativity, spin can also couple to gravitational fields and a gravitomagnetic Stern-Gerlach force arises (see, for example, [24]).
@@ II Notation and Definitions
$x^{\mu}=\begin{pmatrix}ct\\ x\\ y\\ z\end{pmatrix}=\begin{pmatrix}ct\\ \mathbf{r}\end{pmatrix}\ \ \ \ \text{and}\ \ \ \ x_{\mu}=\eta_{\mu\nu}x^{\nu}=\begin{pmatrix}ct\\ -\mathbf{r}\end{pmatrix}\ ,$ (1)
@@
where $\mathbf{v}$ is the three-velocity vector of the particle, and the relativistic factor appearing is defined as $\gamma^{-1}=\sqrt{1-\beta^{2}}$ with $\pmb{\beta}=\frac{\mathbf{v}}{c}$. A Lorentz scalar such as $u_{\mu}u^{\mu}$ is the same in any reference frame, and we are free to evaluate it in the rest frame, where $\mathbf{v}_{\text{rest}}=0$. We see that then $u_{\mu}u^{\mu}=c^{2}$, which holds true in any frame. Taking the derivative with respect to proper time on both sides yields the condition
@@
$\Sigma^{\mu\nu}=\gamma\begin{pmatrix}0&cd_{x}&cd_{y}&cd_{z}\\ -cd_{x}&0&-\mu_{z}&\mu_{y}\\ -cd_{y}&\mu_{z}&0&-\mu_{x}\\ -cd_{z}&-\mu_{y}&\mu_{x}&0\end{pmatrix}\ ,$ (9)
@@
in covariant form, the latter of which is known as the Frenkel condition [3]. For more details on the equivalence of Eqs. (11) and (12), see the Appendix. We need such a constraint to eliminate the degrees of freedom of the tensor that one would otherwise associate with an electric moment independent of the magnetic one. The electromagnetic moment tensor $\Sigma^{\mu\nu}$ has two remaining degrees of freedom – three out of the six initial ones are taken care of by way of Eq. (12) and one is removed by the fixed value of $\mu=|\pmb{\mu}_{\text{rest}}|$ via the Lorentz invariant $\Sigma^{\mu\nu}\Sigma_{\mu\nu}=-2\mu$.
@@ III Evolution of the Electric Sources
where $m$ is the mass and $e$ the elementary charge of the particle. If field gradients are non-negligible, a coupling of the particle trajectory to the magnetic moment arises, i.e., the latter influences the former. Such contributions are given in [29, Eq. (6)], but we do not consider them here. To gain more insight, we rewrite Eq. (13) explicitly in its temporal and spatial components
@@
with the typical transformation parameters – the three angles of rotation represented by $\pmb{\varphi}$ and the rapidity vector $\pmb{\zeta}=\hat{\mathbf{v}}_{b}\tanh^{-1}\frac{v_{b}}{c}$ of the boost with $\hat{\mathbf{v}}_{b}=\frac{\mathbf{v}_{b}}{v_{b}}$ and $v_{b}=\sqrt{\mathbf{v}_{b}\cdot\mathbf{v}_{b}}$ – and their corresponding generators – the vectors of matrices $\mathbf{J}$ and $\mathbf{K}$, respectively, the explicit form of which is given in the Appendix. Applying this Lorentz transformation to the four-velocity $u^{\mu}$ yields the transformed quantity in covariant form and explicitly for the components
@@
$\begin{pmatrix}\gamma^{\prime}c\\ (\gamma\mathbf{v})^{\prime}\end{pmatrix}$ $=\begin{pmatrix}\gamma c\\ \gamma\mathbf{v}\end{pmatrix}+\begin{pmatrix}\pmb{\zeta}\cdot(\gamma\mathbf{v})\\ \gamma c\pmb{\zeta}+\pmb{\varphi}\times(\gamma\mathbf{v})\end{pmatrix}\ .$
@@
$\ \ \ +\Delta\tau\frac{e}{m}\begin{pmatrix}\frac{1}{c}\mathbf{E}\cdot(\gamma\mathbf{v})\\ \gamma c\frac{1}{c}\mathbf{E}-\mathbf{B}\times(\gamma\mathbf{v})\end{pmatrix}$
@@
We can then compare the respective right-hand sides of the explicit forms in Eqs. (16) and (17). The first term – the four-velocity – is the same in both equations. In the respective second terms, very similar expressions are found – in the temporal components of both, there is a scalar product with $\gamma\mathbf{v}$ and in the spatial ones - a term proportional to $\gamma c$ and one consisting of a cross product with $\gamma\mathbf{v}$. We can now directly identify the remaining quantities and make the following substitution of the Lorentz transformation parameters
@@
Like any Lorentz transformation, this one keeps the metric tensor invariant, which is essential for the Lorentz invariance of the infinitesimal four-volume element related to the proper time in Eq. (2). More details regarding the form of the inverse and the proof of the above statements can be found in the Appendix.
@@ IV Evolution of the Magnetic Sources
After analyzing the Lorentz force that drives the electric sources to evolve in the presence of homogeneous electromagnetic fields, we aim to study the behavior of the magnetic sources under such conditions in the following. Although we discuss the electric and magnetic sources in a manner that might suggest that they are separate objects, it is important to note that this is often not the case. For example, all electrically charged fermions, such as the electron, possess an intrinsic half-integer spin, making them both electric and magnetic sources. However, we will treat the dynamics separately and will not assume any a priori coupling between the evolution of the electric sources (the trajectory of the particle) and the magnetic ones (the orientation of the magnetic moment). The only requirement we impose is Eq. (11), or equivalently, Eq. (12), which ensures that there is no EDM in the instantaneous rest frame of the particle ($\mathbf{v}_{\text{rest}}=0$).
@@
$=\Sigma^{\mu\nu}+\Delta\tau\kappa\big{(}F^{\mu}_{\ \alpha}\Sigma^{\alpha\nu}-F^{\nu}_{\ \alpha}\Sigma^{\alpha\mu}\big{)}$
up to first order in $\Delta\tau$. For more details on the above calculation, see the Appendix. Reconstructing the proper time derivative using the same assumption as for the electric sources, ${\Sigma^{\prime}}^{\mu\nu}\rightarrow{\Sigma}^{\mu\nu}(\tau+\Delta\tau)$ and accordingly ${\Sigma}^{\mu\nu}={\Sigma}^{\mu\nu}(\tau)$, in the above equation yields
@@
Importantly, the acquired evolution equation yields a totally antisymmetric tensor on both hand sides. It is written in three different but equivalent forms that we can compare to the evolution equations for classical intrinsic spin found in literature, that were obtained by different means. The expressions in the three lines in Eq. (23) correspond to those in [10], [11] and [9], respectively. Comparison with the evolution equations in [10, 11, 16] yields the explicit expression for the prefactor $\kappa$  [^1] The given $\kappa$ has been adjusted from the one in Refs. [10, 11], which have an additional factor of $c$ in the denominator due to a difference in the definitions of the electromagnetic field and moment tensor elements.
@@
$\rightarrow U^{\prime}$ $=-\frac{1}{2}\Sigma_{\mu\nu}(F^{\mu\nu}+K^{\mu}u^{\nu}-K^{\nu}u^{\mu})$
$=-\frac{1}{2}\Sigma_{\mu\nu}{F^{\prime}}^{\mu\nu}\ .$
We then calculate the equation of motion anew by substituting $F^{\mu\nu}$ with ${F^{\prime}}^{\mu\nu}$, which is now an effective field containing the auxiliary four-vector $K^{\mu}$. The added terms in the energy functional vanish via the Frenkel condition, Eq. (12). We do not make any assumptions about $K^{\mu}$ and instead let its form emerge naturally. Such a derivation for an electron was partially done already by Frenkel [3], but the missing knowledge at the time that the electron $g$-factor is, in fact, $g\neq 2$ yielded incomplete results. We take a general gyromagnetic ratio $g$ into account here. Replacing $F^{\mu\nu}$ by ${F^{\prime}}^{\mu\nu}$ in the commutator equation, Eq. (23) leads to
@@
$\ \ \ \ -\frac{eg}{2m}K_{\alpha}\big{(}u^{\mu}\Sigma^{\alpha\nu}-u^{\nu}\Sigma^{\alpha\mu}\big{)}$
and we can once again use the Frenkel condition to find the unknown four-vector $K^{\mu}$
$K^{\mu}=\frac{1}{gc^{2}}(g-2)F^{\nu\mu}u_{\nu}\ .$ (32)
The detailed derivation can be found in the Appendix. We note that $K^{\mu}$ is space-like, i.e., $K^{\mu}u_{\mu}=0$. The evolution equation for the electromagnetic moment tensor hence becomes
@@
The above equation agrees with the evolution equations in [10, 11], which are derived in different ways than the one above. We have now found an additional term, which is proportional to $(g-2)$ and corresponds to the contribution of an anomalous magnetic moment that an electrically charged fermion might have. We know that the intrinsic magnetic moments of the electron and the muon do have an anomalous part [35]. We see that in this classical treatment, the corresponding contribution arises naturally when we keep $g$ general and do not set it to $2$. As we see in the second term on the right-hand side of Eq. (33), the evolution of the moment tensor is coupled to the four-velocity of the particle, i.e., its trajectory, through the electromagnetic field. This was not the case in Eq. (23), where $\Sigma^{\mu\nu}$ evolves independently of $u^{\mu}$. For $g=2$ in Eq. (33), we recover Eq. (23).
@@
Furthermore, we take ${F^{\prime}}^{\mu\nu}$ and write a Lorentz-like transformation with it in analogy to the case with just $F^{\mu\nu}$
${\Lambda^{\prime}}^{\mu}_{\ \nu}=\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}{F^{\prime}}^{\mu}_{\ \nu}$ (34)
@@
and call it an effective Lorentz-like transformation. Transforming the electromagnetic moment tensor according to the transformation rules used in Eq. (22) and reconstructing the proper time derivative by utilizing the method that led to Eq. (23), now yields Eq. (33). Hence, we can keep the idea of the electromagnetic fields driving the sources by means of Lorentz boosts and rotations. In this case, we have an effective electromagnetic tensor ${F^{\prime}}^{\mu\nu}$, or adjusted boost and rotation parameters, which act as a coupler between the evolution of the magnetic sources and the trajectory of the particle.
@@
The effective Lorentz-like transformation in Eq. (34) originates from adjusting the Lorentz-like transformation acting on the electric sources, Eq. (20), to fit the properties and constraints of the evolution of magnetic sources, leading to Eq. (33). However, we can show that when we let the effective Lorentz-like transformation act on the electric sources, that is, the four-velocity $u^{\mu}$, we reproduce the Lorentz force again. The reason why both $\frac{e}{m}{F}^{\mu}_{\ \nu}$ and $\frac{eg}{2m}{F^{\prime}}^{\mu}_{\ \nu}$ in Eqs. (20) and (34), respectively, yield the correct result for the electric sources is the following. In the latter case the emerging terms proportional to $g$ cancel, while the remaining term yields the prefactor $\frac{e}{m}$ in the Lorentz force, which the former expression does directly. The explicit calculation is given in the Appendix. Importantly, we can now use a single transformation, namely, the effective Lorentz-like one given in Eq. (34), apply it to both fundamental sources – electric and magnetic – and correctly recreate their respective evolution equations in the presence of electromagnetic fields, which is a quite satisfactory unification. Hence, we can interpret the driving forces of both electric and magnetic sources as a series of infinitesimal boosts and rotations containing effective electromagnetic fields.

### v2

@@ I Introduction and Summary
For particles with spin in very strong or inhomogeneous fields additional terms arise, which originate from a Stern-Gerlach-type force. A covariant formulation of the Stern-Gerlach force was first considered by Frenkel [3] and then by Good [26] and Nyborg [27], among others. It has been revisited and extended in recent years [28], and the dynamics of both neutral and charged particles were studied in [29, 30, 31]. We do not consider such contributions here since we focus on homogeneous fields and fields, whose gradients are negligibly small. Further extensions to the dynamical equations and implications thereof are presented in [32]. In the framework of general relativity, spin can also couple to gravitational fields and a gravitomagnetic Stern-Gerlach force arises (see, for example, [24]).
@@ II Notation and Definitions
$x^{\mu}=\begin{pmatrix}ct\\ x\\ y\\ z\end{pmatrix}=\begin{pmatrix}ct\\ \mathbf{r}\end{pmatrix}\ \ \ \ \text{and}\ \ \ \ x_{\mu}=\eta_{\mu\nu}x^{\nu}=\begin{pmatrix}ct\\ -\mathbf{r}\end{pmatrix}\$, (1)
@@
where $\mathbf{v}$ is the three-velocity vector of the particle, and the relativistic factor appearing is defined as $\gamma^{-1}=\sqrt{1-\beta^{2}}$ with $\pmb{\beta}=\frac{\mathbf{v}}{c}$. A Lorentz scalar is a constant in any reference frame, and in the case of the projection of the four-velocity onto itself, that value is $u_{\mu}u^{\mu}=c^{2}$. Taking the derivative with respect to proper time on both sides yields the condition
@@
$\Sigma^{\mu\nu}=\gamma\begin{pmatrix}0&cd_{x}&cd_{y}&cd_{z}\\ -cd_{x}&0&-\mu_{z}&\mu_{y}\\ -cd_{y}&\mu_{z}&0&-\mu_{x}\\ -cd_{z}&-\mu_{y}&\mu_{x}&0\end{pmatrix}\$, (9)
@@
in covariant form, the latter of which is known as the Frenkel condition [3]. For more details on the equivalence of Eqs. (11) and (12), see Appendix A. We need such a constraint to eliminate the degrees of freedom of the tensor that one would otherwise associate with an electric moment independent of the magnetic one. The electromagnetic moment tensor $\Sigma^{\mu\nu}$ has two remaining degrees of freedom – three out of the six initial ones are taken care of by way of Eq. (12) and one is removed by the fixed value of $\mu=|\pmb{\mu}_{\text{rest}}|$ via the Lorentz invariant $\Sigma^{\mu\nu}\Sigma_{\mu\nu}=-2\mu$.
@@ III Evolution of the Electric Sources
where $m$ is the mass and $e$ the elementary charge of the particle. If field gradients are non-negligible, a coupling of the particle trajectory to the magnetic moment arises, i.e., the latter influences the former. Such contributions are given in [29, Eq. (6)], but we do not consider them here. It is worth noting that in the setting of homogeneous fields, we can treat the field tensor as just another constant and solve the ordinary differential equation for $u^{\mu}$ given by the Lorentz force, Eq. (13). The expression for the trajectory then contains an exponentiation of the proper time and the electromagnetic field tensor (see [35, Chapter 1.1.3]). To gain more insight, we rewrite Eq. (13) explicitly in its temporal and spatial components
@@
with the typical transformation parameters – the three angles of rotation represented by $\pmb{\varphi}$ and the rapidity vector $\pmb{\zeta}=\hat{\mathbf{v}}_{b}\tanh^{-1}\frac{v_{b}}{c}$ of the boost with $\hat{\mathbf{v}}_{b}=\frac{\mathbf{v}_{b}}{v_{b}}$ and $v_{b}=\sqrt{\mathbf{v}_{b}\cdot\mathbf{v}_{b}}$ – and their corresponding generators – the vectors of matrices $\mathbf{J}$ and $\mathbf{K}$, respectively, the explicit form of which is given in Appendix B.1. Applying this Lorentz transformation to the four-velocity $u^{\mu}$ yields the transformed quantity in covariant form and explicitly for the components
@@
$\begin{pmatrix}\gamma^{\prime}c\\ (\gamma\mathbf{v})^{\prime}\end{pmatrix}$ $=\begin{pmatrix}\gamma c\\ \gamma\mathbf{v}\end{pmatrix}+\begin{pmatrix}\pmb{\zeta}\cdot(\gamma\mathbf{v})\\ \gamma c\pmb{\zeta}+\pmb{\varphi}\times(\gamma\mathbf{v})\end{pmatrix}\$.
@@
$\ \ \ +\Delta\tau\frac{e}{m}\begin{pmatrix}\frac{1}{c}\mathbf{E}\cdot(\gamma\mathbf{v})\\ \gamma c\frac{1}{c}\mathbf{E}-\mathbf{B}\times(\gamma\mathbf{v})\end{pmatrix}\$
@@
We can then compare the respective right-hand sides of the explicit forms in Eqs. (16) and (17). The first term – the four-velocity – is the same in both equations. In the respective second terms, very similar expressions are found; in the temporal components of both, there is a scalar product with $\gamma\mathbf{v}$ and in the spatial ones – a term proportional to $\gamma c$ and one consisting of a cross product with $\gamma\mathbf{v}$. We can now directly identify the remaining quantities and make the following substitution of the Lorentz transformation parameters
@@
Like any Lorentz transformation, this one keeps the metric tensor invariant, which is essential for the Lorentz invariance of the infinitesimal four-volume element related to the proper time in Eq. (2). More details regarding the form of the inverse and the proof of the above statements can be found in Appendices B.2 and B.3.
@@ IV Evolution of the Magnetic Sources
After analyzing the Lorentz force that drives the electric sources to evolve in the presence of homogeneous electromagnetic fields, we aim to study the behavior of the magnetic sources under such conditions in the following. Although we discuss the electric and magnetic sources in a manner that might suggest that they are separate objects, it is important to note that this is often not the case. For example, all electrically charged fermions, such as the electron, possess an intrinsic half-integer spin, making them both electric and magnetic sources simultaneously. However, since we only consider the dynamics of the sources in homogeneous fields, we will treat it separately and will not assume any a priori coupling between the evolution of the electric sources (the trajectory of the particle) and the magnetic ones (the orientation of the magnetic moment) [^1] In the case of fields, whose gradients are non-negligible, a coupling between the electric and magnetic source dynamics arises (see [29, Eq. (6)]) and we would not be allowed to assume independence here.. The only requirement we impose is Eq. (11), or equivalently, Eq. (12), which ensures that there is no EDM in the instantaneous rest frame of the particle ($\mathbf{v}_{\text{rest}}=0$).
@@
$=\Sigma^{\mu\nu}+\Delta\tau\kappa\big{(}F^{\mu}_{\ \alpha}\Sigma^{\alpha\nu}-F^{\nu}_{\ \alpha}\Sigma^{\alpha\mu}\big{)}\$
up to first order in $\Delta\tau$. For more details on the above calculation, see Appendix C.1. Reconstructing the proper time derivative using the same assumption as for the electric sources, ${\Sigma^{\prime}}^{\mu\nu}\rightarrow{\Sigma}^{\mu\nu}(\tau+\Delta\tau)$ and accordingly ${\Sigma}^{\mu\nu}={\Sigma}^{\mu\nu}(\tau)$, in the above equation yields
@@
Importantly, the acquired evolution equation yields a totally antisymmetric tensor on both hand sides. It is written in three different but equivalent forms that we can compare to the evolution equations for classical intrinsic spin found in literature, which were obtained by different means. The expressions in the three lines in Eq. (23) correspond to those in [10], [11] and [9], respectively. Comparison with the evolution equations in [10, 11, 16] yields the explicit expression for the prefactor $\kappa$  [^2] The given $\kappa$ has been adjusted from the one in Refs. [10, 11], which have an additional factor of $c$ in the denominator due to a difference in the definitions of the electromagnetic field and moment tensor elements.
@@
$\rightarrow\tilde{U}$ $=-\frac{1}{2}\Sigma_{\mu\nu}(F^{\mu\nu}+\Xi^{\mu}u^{\nu}-\Xi^{\nu}u^{\mu})$
$=-\frac{1}{2}\Sigma_{\mu\nu}{\tilde{F}}^{\mu\nu}\ .$
We then calculate the equation of motion anew by substituting $F^{\mu\nu}$ with ${\tilde{F}}^{\mu\nu}$, which is now an effective field containing the auxiliary four-vector $\Xi^{\mu}$. The added terms in the energy functional vanish via the Frenkel condition, Eq. (12). We do not make any assumptions about $\Xi^{\mu}$ and instead let its form emerge naturally. Such a derivation for an electron was partially done already by Frenkel [3], but the missing knowledge at the time that the electron $g$-factor is, in fact, $g\neq 2$ yielded incomplete results. We take a general gyromagnetic ratio $g$ into account here. Replacing $F^{\mu\nu}$ by ${\tilde{F}}^{\mu\nu}$ in the commutator equation, Eq. (23) leads to
@@
$\ \ \ \ -\frac{eg}{2m}\Xi_{\alpha}\big{(}u^{\mu}\Sigma^{\alpha\nu}-u^{\nu}\Sigma^{\alpha\mu}\big{)}\$
and we can once again use the Frenkel condition to find the unknown four-vector $\Xi^{\mu}$
$\Xi^{\mu}=\frac{1}{gc^{2}}(g-2)F^{\nu\mu}u_{\nu}\ .$ (32)
The detailed derivation can be found in Appendix C.2. We note that $\Xi^{\mu}$ is space-like, i.e., $\Xi^{\mu}u_{\mu}=0$. The evolution equation for the electromagnetic moment tensor hence becomes
@@
The above equation agrees with the evolution equations in [10, 11], which are derived in different ways than the one above. We have now found an additional term, which is proportional to (g2)(g-2) and corresponds to the contribution of an anomalous magnetic moment that an electrically charged fermion might have. We know that the intrinsic magnetic moments of the electron and the muon do have an anomalous part [36]. We see that in this classical treatment, the corresponding contribution arises naturally when we keep $g$ general and do not set it to $2$. As we see in the second term on the right-hand side of Eq. (33), the evolution of the moment tensor is coupled to the four-velocity of the particle, i.e., its trajectory, through the electromagnetic field. This was not the case in Eq. (23), where $\Sigma^{\mu\nu}$ evolves independently of $u^{\mu}$. For $g=2$ in Eq. (33), we recover Eq. (23).
@@
Furthermore, we take the effective electromagnetic field tensor ${\tilde{F}}^{\mu\nu}$ and write a Lorentz-like transformation with it in analogy to the case with just $F^{\mu\nu}$
${\tilde{\Lambda}}^{\mu}_{\ \nu}$ $=\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}\begin{pmatrix}0&\tilde{E}_{x}/c&\tilde{E}_{y}/c&\tilde{E}_{z}/c\\ \tilde{E}_{x}/c&0&\tilde{B}_{z}&-\tilde{B}_{y}\\ \tilde{E}_{y}/c&-\tilde{B}_{z}&0&\tilde{B}_{x}\\ \tilde{E}_{z}/c&\tilde{B}_{y}&-\tilde{B}_{x}&0\end{pmatrix}$ (34)
$=\delta^{\mu}_{\nu}+\Delta\tau\frac{eg}{2m}{\tilde{F}}^{\mu}_{\ \nu}$
with the effective fields explicitly given by
$\tilde{\mathbf{E}}$ $=\mathbf{E}+\frac{(g-2)}{gc^{2}}\gamma^{2}\big{(}\mathbf{v}(\mathbf{v}\cdot\mathbf{E})-c^{2}(\mathbf{E}+\mathbf{v}\times\mathbf{B})\big{)}$ (35)
$\tilde{\mathbf{B}}$ $=\mathbf{B}-\frac{(g-2)}{gc^{2}}\gamma^{2}\big{(}\mathbf{v}(\mathbf{v}\cdot\mathbf{B})-v^{2}\mathbf{B}+\mathbf{v}\times\mathbf{E}\big{)}$
and call it an effective Lorentz-like transformation. Analogously to ${\Lambda}^{\mu}_{\ \nu}$, its inverse is given by
${\tilde{\Lambda}}_{\nu}^{\ \mu}=({\tilde{\Lambda}}^{-1})^{\mu}_{\ \nu}=\delta^{\mu}_{\nu}-\Delta\tau\frac{eg}{2m}{\tilde{F}}^{\mu}_{\ \nu}\ .$ (36)
and it keeps the metric tensor invariant. This can be readily seen from calculations equivalent to those in Appendix B.2 and B.3 with the substitution $\frac{e}{m}{F}^{\mu}_{\ \nu}\rightarrow\frac{eg}{2m}{\tilde{F}}^{\mu}_{\ \nu}$. Transforming the electromagnetic moment tensor according to the transformation rules used in Eq. (22) and reconstructing the proper time derivative by utilizing the method that led to Eq. (23), now yields Eq. (33). The effective Lorentz-like transformation can still be expressed in terms of the Lorentz boost and rotation generators, $\mathbf{K}$ and $\mathbf{J}$, respectively, as in Eq. (15), with parameters
$\pmb{\zeta}\rightarrow\Delta\tau\frac{eg}{2m}\frac{\tilde{\mathbf{E}}}{c}\ ,\ \pmb{\varphi}\rightarrow-\Delta\tau\frac{eg}{2m}\tilde{\mathbf{B}}\$. (37)
@@
with $\tilde{\mathbf{E}}$ and $\tilde{\mathbf{B}}$ given in Eq. (35). Hence, we can keep the idea of the electromagnetic fields driving the sources by means of Lorentz boosts and rotations. In this case, we have a totally antisymmetric effective electromagnetic tensor ${\tilde{F}}^{\mu\nu}$, or adjusted boost and rotation parameters, which act as a coupler between the evolution of the magnetic sources and the trajectory of the particle.
@@
The effective Lorentz-like transformation in Eq. (34) originates from adjusting the Lorentz-like transformation acting on the electric sources, Eq. (20), to fit the properties and constraints of the evolution of magnetic sources, leading to Eq. (33). However, we can show that when we let the effective Lorentz-like transformation act on the electric sources, that is, the four-velocity $u^{\mu}$, we reproduce the Lorentz force again. The reason why both $\frac{e}{m}{F}^{\mu}_{\ \nu}$ and $\frac{eg}{2m}{\tilde{F}}^{\mu}_{\ \nu}$ in Eqs. (20) and (34), respectively, yield the correct result for the electric sources is the following. In the latter case the emerging terms proportional to $g$ cancel, while the remaining term yields the prefactor $\frac{e}{m}$ in the Lorentz force, which the former expression does directly. The explicit calculation is given in Appendix C.3. As the arising differential equation remains the same, Eq. (13), the exponentiation given in [35, Chapter 1.1.3] is still its solution. Importantly, we can now use a single transformation, namely, the effective Lorentz-like one given in Eq. (34), apply it to both fundamental sources – electric and magnetic – and correctly recreate their respective evolution equations in the presence of electromagnetic fields, which is a quite satisfactory unification. Hence, we can interpret the driving forces of both electric and magnetic sources as a series of infinitesimal boosts and rotations containing effective electromagnetic fields.